you assume that Coca-Cola company is the representative firm & has a total cost given by TC= 125+q2+q where q is the quantity of the output that the firm produces. You further assume that the market demand for Coca-Cola products is given by the equation P= 1250.5-2Q, where Q is the entire market quantity. In addition, you are informed that the supply curve for the market is given by P=150+Q.
a) What is the equilibrium quantity and price in this market given this information?
b) Given the answer in part (a) above, obtain the profit maximizing level of output, the total revenue, total cost and profit at the market equilibrium.
Is this a short run or long run equilibrium? Explain your answer.
c) Given your answer in (b) above, what do u think will happen in this market in the long run?
1
Expert's answer
2017-09-12T02:27:06-0400
a) Equilibrium quantity:
1250.5-2Q=150+Q
3Q=1100.5
Q=367
Equilibrium price:
P=150+367=517
b) TR = P*Q = 517*367 = 189739
TC = 125 + 367^2+367 = 135181
PR = TR-TC = 189739-135181 = 54558
ATC = 125/Q+Q = 368
It is a short run, because P>ATC
c) It the long run in this market there may be surplus.